OpenStudy (anonymous):

HELP PLEASE Complete the table for the function and find the indicated limit. A. -1.22843; -1.20298; -1.20030; -1.19970; -1.19699; -1.16858 limit = -1.20 B. -2.18529; -2.10895; -2.10090; -2.09910; -2.09096; -2.00574 limit = -2.10 C. -4.09476; -4.00995; -4.00100; -3.99900; -3.98995; -3.89526 limit = -4.0 D. 4.09476; 4.00995; 4.00100; 3.99900; 3.98995; 3.89526 limit = 4.0

OpenStudy (hunus):

What's the function?

OpenStudy (hunus):

Plug in the values you're given in the table into the function and see what's equal to for those values

OpenStudy (anonymous):

i GOT EITHER C OR B

OpenStudy (anonymous):

I MEAN C OR D

OpenStudy (hunus):

For the first value, x = -0.1, we get \[\frac{(-0.1)^3-6(-0.1)+8}{(-0.1)-2}=\frac{-0.001+0.6+8}{-2.1}=\frac{8.599}{-2.1}=-4.094\]

OpenStudy (anonymous):

SO IM CORRCT ITS C

OpenStudy (hunus):

Yes. What happens when you actually plug x=0 into the equation?

OpenStudy (anonymous):

YOU GET YOUR ANSWER...THANKS @Hunus

OpenStudy (hunus):

Yup. You can't always just plug in the value though. You can here because the graph is actually continuous at the point x = 0.

OpenStudy (anonymous):

THANKS AGAIN @Hunus ;-)